OpenStudy (anonymous):

a facuet is used to add a large bottle that already contained some water. After it has been filling for 3 seconds the gauge on the bottle indicates that it contains 16 ounces of water. After it has been filling for 10 seconds, the gauge indicates the bottle contains 44 ounces of water. Let y be the amount of water in the bottle x seconds after the facuet was turned on. Write a linear equation that relates the amount of water in the bottle y to the time x

OpenStudy (anonymous):

\[\frac{ \Delta V }{ \Delta t }\] The change in volume, divided by the change in time will you give the slope of a linear equation, then you just have to solve for the constant!

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

Okay, a linear equation is in this format \[y = mx + b\] where m is the slope, and b is a constant.

OpenStudy (anonymous):

still confused on the slope and constant

OpenStudy (anonymous):

Okay, slope (m) is the change in y divided by the change in x (rise over run), the constant is the y intercept, which is what y will equal when x is 0

OpenStudy (anonymous):

ok but in m problem what is that

OpenStudy (anonymous):

Okay so \[m = \frac{ \Delta V }{ \Delta t } = \frac{ 44oz - 16oz }{ 10s - 3s }\]

OpenStudy (anonymous):

so i get 28/7=4 which means than y=4x+4

OpenStudy (anonymous):

Yes! Exactly! Great job!